# Connected Components in Linear Work and Near-Optimal Time

### v2

@@ 6 Stage 3: Connectivity on the Sampled Graph
This assumption will be removed in §7 to give a connected component algorithm for any graphs with any spectral gap and the algorithm does not need any prior knowledge on the spectral gap.
@@
SampleSolve$(G=(V,E))$: 1. If $|V|\leq n^{0.1}$ then removes parallel edges and loops from $E$ and compute the connected components of graph $(V,E)$ by Theorem 2 and goto Step 4. 2. Sample each edge in $E$ w.p. $1/(\log n)^{7}$ to obtain graph $G^{\prime}$. 3. Compute the connected components of graph $G^{\prime}$ by Theorem 2. 4. For each vertex $v$ in the original input graph: $v.p=v.p.p.p$.
@@ 7.1 Algorithmic Framework
Connectivity$(G=(V,E))$: 1. Initialize the labeled digraph: for each vertex $v\in V$: $v.p=v$. 2. Run Reduce$(V,E,k=10^{6}\log\log n)$ to obtain the current graph $G^{\prime}$. 3. Sample each edge of $G^{\prime}$ w.p. $1/(\log n)^{7}$ independently to get random subgraph $H_{1}$. Sample each edge of $G^{\prime}$ w.p. $1/(\log n)^{7}$ independently to get random subgraph $H_{2}$. The sampled graphs are copied out so that $G^{\prime}$ is unchanged. 4. Copy the edges of $G^{\prime}$ to create an edge set $E_{\text{filter}}$. 5. For phase $i$ from 00 to $10\log\log n$: $E_{\text{filter}}=\textsc{Interweave}(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$, if $E_{\text{filter}}=\emptyset$ then goto Step 6. 6. Shortcut$(V)$.
@@
Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$: 1. Set $b=(\log n)^{100\cdot 1.1^{i}}$. 2. $\textsc{Increase}(G^{\prime},H_{1},H_{2},b)$ — see § 7.3 for more details. 3. Run $20\log b$ rounds of $H_{1}=\textsc{Expand-Maxlink}(H_{1})$ then run $10^{4}\log\log n$ rounds of the algorithm in Theorem 2 on $H_{1}$. Alter$(H_{1})$. 4. If all edges in $E(H_{1})$ are loops then call Remain$(G^{\prime},H_{1})$ and return $\emptyset$. 5. Revert the labeled digraph and graph $H_{1}$ to their states in Step 1. 6. Repeat for $10^{6}\cdot 1.1^{i}\log\log n$ rounds: Matching$(E_{\text{filter}})$, Alter$(E_{\text{filter}})$, delete each edge from $E_{\text{filter}}$ w.p. $10^{-4}$. 7. Repeat for $i+2\log\log n$ times: Shortcut$(V(G^{\prime}))$. 8. Edge set $E^{\prime}=\{(v,w)\in E(G^{\prime})\mid v.p\in V(G^{\prime})\backslash V(E_{\text{filter}})\}$. Alter$(E^{\prime})$. 9. Repeat for $10^{6}\cdot 1.1^{i}\log\log n$ rounds: Matching$(E^{\prime})$, Shortcut$(V(G^{\prime}))$, Alter$(E^{\prime})$. 10. Reverse$(V(E_{\text{filter}}),E(H_{2}))$. 11. Return $E_{\text{filter}}$.
Remain$(G^{\prime},H_{1})$: 1. Alter$(E(G^{\prime}))$. 2. Edge set $E_{\text{remain}}=E(G^{\prime})\backslash E(H_{1})$. 3. Remove loops and parallel edges from $E_{\text{remain}}$. 4. Compute the connected components of graph $(V(E_{\text{remain}}),E_{\text{remain}})$ by Theorem 2.
@@ 7.3 Reducing the Work of Increase
Increase$(G^{\prime},H_{1},H_{2},b)$: 1. Edge set $E_{H}=\textsc{SparseBuild}(G^{\prime},H_{2},b)$, vertex set $V=V(G^{\prime})$, graph $H=(V(E_{H}),E_{H})$. 2. Run Step 2 to Step 9 in the Increase$(V,E,b)$ in §5.3. 3. Alter$(E(H_{1}))$.
@@ 7.3.2 Increasing Vertex Degrees by Sparse Skeleton Graph
[…]
@@ 7.4.1 Building the Auxiliary Array
BuildAuxiliary$(G^{\prime})$: 1. Padded sort all edges of $G^{\prime}$ by their first ends and store the output array $\mathcal{A}$. 2. For each item $(v,w)\in\mathcal{A}$ in parallel: if the first end of the predecessor of $(v,w)$ in $\mathcal{A}$ is not $v$ then write the index of $(v,w)$ in $\mathcal{A}^{\prime}$ into $v.l$ in the private memory of (the processor corresponding to) $v$. 3. For each item $(v,w)\in\mathcal{A}$ in parallel: if the first end of the successor of $(v,w)$ in $\mathcal{A}$ is not $v$ then write the index of $(v,w)$ in $\mathcal{A}$ into $v.r$ in the private memory of $v$. 4. For each vertex $v\in V(G^{\prime})$: if $v.r-v.l\geq(\log n)^{90}$ then (approximate) compact the subarray from index $v.l$ to index $v.r$ (inclusive) in $\mathcal{A}$.
@@ 7.4.3 Computing the Edge Set in Interweave
[…]
@@
[…]
@@
[…]
@@ 7.6 The Running Time and Correctness of Connectivity
[…]
@@
[…]
@@ 8.1 Bounding the Size of Edge Set with High Probability
[…]
@@
[…]
@@ Appendix A Conditional Lower Bound
In this section, we show that our time bound is optimal when the input graph has component-wise spectral gap at most $1/\operatorname{poly}(\log(n))$ conditioning on the widely believed 2-Cycle Conjecture (see below). The approach is similar to that of [${}^{+}$19], who gives an $\Omega(\log d)$ conditional lower bound, where $d\geq\operatorname{poly}(\log(n))$ is the diameter.
Conjecture A.1 (2-Cycle Conjecture [RVW18, ASW19, ${}^{+}$18, ${}^{+}$19]).
@@
For simplicity, we state the theorem in the MPC model so that our proof can be based on the MPC model. The conditional lower bound easily holds in any PRAM model with $O(m+n)$ total memory by a simple simulation [ASW19, ${}^{+}$18]: if there exists a PRAM algorithm that distinguishes between one cycle from two cycles in $o(\log n)$ time, then an MPC with $n^{1-\Omega(1)}$ space per processor can simulate this PRAM algorithm with $O(1)$ overhead, which violates the conjecture.
@@
Each round takes $o(\log(1/\lambda^{\prime}))$ rounds w.h.p.
@@
The edge deletion at the beginning of the round and the contraction at the end of the round can be implemented in $O(1)$ MPC rounds (e.g., see [${}^{+}$18]). It remains to show that the graph $G^{\prime}$ ($G$ after edge deletions) has spectral gap at least $\lambda^{\prime}$ w.h.p. at the beginning of each round. [^22] If graph $G^{\prime}$ has component-wise spectral gap greater than $\lambda^{\prime}$, the algorithm creates a dummy component of spectral gap $\lambda^{\prime}$ and adds it into $G^{\prime}$ such that $G^{\prime}$ has component-wise spectral gap exactly $\lambda^{\prime}$ and thus ALG is applicable on $G^{\prime}$. The dummy component has no edge with the original $G^{\prime}$ so does not influence the computation of other components. Moreover, the dummy component has at most $m$ edges (e.g. a path of fixed length) and can be created and deleted at the end of this round in $O(1)$ MPC rounds.
@@
By Lemma A.3, the total number of round is at most
@@ Appendix B Edge Sampling Does Not Preserve Diameter
In this section we show that edge sampling does not preserve the diameter of graphs. We provide an undirected graph $G$ with a diameter of $O(\log n)$ such that if we sample every edge in $G$ with probability of $p=\frac{1}{\log n}$, the largest diameter in the connected components of $G$ would be at least $\frac{n}{\operatorname{poly}(\log n)}$ where $n$ is the number of vertices in the graph. This shows that edge sampling can dramatically increase the diameter of a graph.
@@
To show the connectivity, first consider the vertices $z_{1},\cdots,z_{(\log n)^{3}}$. We show that these vertices form a connected subgraph in $\widetilde{G}$. Notice that these vertices form a complete induced subgraph in $G$. Since we sample each edge with probability $p=1/\log n$, we can say that in $\widetilde{G}$, the subgraph induced by these vertices has the same distribution as $G((\log n)^{3},p)$. Here, $G(n,p)$ is used to denote the Erdős–Rényi graph with $n$ vertices where each edge is included in it independently with probability $p$. It is known that $G(n,p)$ is connected with probability $e^{-e^{-c}}$ if $p\geq\frac{(1+c)\ln n}{n}$ [${}^{+}$60]. Recall that in the random subgraph $G((\log n)^{3},p)$, we have $p=1/\log n$. Thus, $p\gg\frac{3\ln n\ln((\log n)^{3})}{(\log n)^{3}}$ for large enough $n$ which implies the connectivity of this random graph w.p. $1-1/n^{2}$.

### v4

@@ 6 Stage 3: Connectivity on the Sampled Graph
This assumption will be removed in §7 to give a connected components algorithm for any graphs with any spectral gap and the algorithm does not need any prior knowledge on the spectral gap.
@@
SampleSolve$(G=(V,E))$:
  1. 1.
    If $|V|\leq n^{0.1}$ then removes parallel edges and loops from $E$ and compute the connected components of graph $(V,E)$ by Theorem 2 and goto Step 4.
  2. 2.
    Sample each edge in $E$ w.p. $1/(\log n)^{7}$ to obtain graph $G^{\prime}$.
  3. 3.
    Compute the connected components of graph $G^{\prime}$ by Theorem 2.
  4. 4.
    For each vertex $v$ in the original input graph: $v.p=v.p.p.p$.
@@ 7.1 Algorithmic Framework
Connectivity$(G=(V,E))$:
  1. 1.
    Initialize the labeled digraph: for each vertex $v\in V$: $v.p=v$.
  2. 2.
    Run Reduce$(V,E,k=10^{6}\log\log n)$ to obtain the current graph $G^{\prime}$.
  3. 3.
    Sample each edge of $G^{\prime}$ w.p. $1/(\log n)^{7}$ independently to get random subgraph $H_{1}$. Sample each edge of $G^{\prime}$ w.p. $1/(\log n)^{7}$ independently to get random subgraph $H_{2}$. The sampled graphs are copied out so that $G^{\prime}$ is unchanged.
  4. 4.
    Copy the edges of $G^{\prime}$ to create an edge set $E_{\text{filter}}$.
  5. 5.
    For phase $i$ from 00 to $10\log\log n$: $E_{\text{filter}}=\textsc{Interweave}(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$, if $E_{\text{filter}}=\emptyset$ then goto Step 6.
  6. 6.
    Shortcut$(V)$.
@@
Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$:
  1. 1.
    Set $b=(\log n)^{100\cdot 1.1^{i}}$.
  2. 2.
    $\textsc{Increase}(G^{\prime},H_{1},H_{2},b)$ — see § 7.3 for more details.
  3. 3.
    Run $20\log b$ rounds of $H_{1}=\textsc{Expand-Maxlink}(H_{1})$ then run $10^{4}\log\log n$ rounds of the algorithm in Theorem 2 on $H_{1}$. Alter$(H_{1})$.
  4. 4.
    If all edges in $E(H_{1})$ are loops then call Remain$(G^{\prime},H_{1})$ and return $\emptyset$.
  5. 5.
    Revert the labeled digraph and graph $H_{1}$ to their states in Step 1.
  6. 6.
    Repeat for $10^{6}\cdot 1.1^{i}\log\log n$ rounds: Matching$(E_{\text{filter}})$, Alter$(E_{\text{filter}})$, delete each edge from $E_{\text{filter}}$ w.p. $10^{-4}$.
  7. 7.
    Repeat for $i+2\log\log n$ times: Shortcut$(V(G^{\prime}))$.
  8. 8.
    Edge set $E^{\prime}=\{(v,w)\in E(G^{\prime})\mid v.p\in V(G^{\prime})\backslash V(E_{\text{filter}})\}$. Alter$(E^{\prime})$.
  9. 9.
    Repeat for $10^{6}\cdot 1.1^{i}\log\log n$ rounds: Matching$(E^{\prime})$, Shortcut$(V(G^{\prime}))$, Alter$(E^{\prime})$.
  10. 10.
    Reverse$(V(E_{\text{filter}}),E(H_{2}))$.
  11. 11.
    Return $E_{\text{filter}}$.
Remain$(G^{\prime},H_{1})$:
  1. 1.
    Alter$(E(G^{\prime}))$.
  2. 2.
    Edge set $E_{\text{remain}}=E(G^{\prime})\backslash E(H_{1})$.
  3. 3.
    Remove loops and parallel edges from $E_{\text{remain}}$.
  4. 4.
    Compute the connected components of graph $(V(E_{\text{remain}}),E_{\text{remain}})$ by Theorem 2.
@@ 7.3 Reducing the Work of Increase
Increase$(G^{\prime},H_{1},H_{2},b)$:
  1. 1.
    Edge set $E_{H}=\textsc{SparseBuild}(G^{\prime},H_{2},b)$, vertex set $V=V(G^{\prime})$, graph $H=(V(E_{H}),E_{H})$.
  2. 2.
    Run Step 2 to Step 9 in the Increase$(V,E,b)$ in §5.3.
  3. 3.
    Alter$(E(H_{1}))$.
@@ 7.3.1 Building Sparse Skeleton Graph
SparseBuild$(G^{\prime},H_{2},b)$:
  1. 1.
    For each active root $v\in V(G^{\prime})$: assign a block of $b^{9}$ processors, which will be used as an indexed hash table $\mathcal{H}(v)$ of size $b^{9}$.
  2. 2.
    For each edge $(u,v)\in E(H_{2})$: write $(u,v)$ into a random indexed cell of $\mathcal{H}(v)$.
  3. 3.
    For each active root $v\in V(G^{\prime})$: if the number of items in $\mathcal{H}(v)$ is more than $b^{8}$ then mark $v$ as high, else mark $v$ as low.
  4. 4.
@@ 7.3.2 Increasing Vertex Degrees by Sparse Skeleton Graph
During the execution of Connectivity$(G)$, w.p. $1-1/(\log n)^{5}$, for any integer $i\in[0,10\log\log n)$ and any active root $v\in V(G^{\prime})$ at the end of Step 2 within Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$, it must be $\overline{\deg}_{G^{\prime}}(v)\geq b=(\log n)^{100\cdot 1.1^{i}}$.
@@ 7.4.1 Building the Auxiliary Array
BuildAuxiliary$(G^{\prime})$:
  1. 1.
    Padded sort all edges of $G^{\prime}$ by their first ends and store the output array $\mathcal{A}$.
  2. 2.
    For each item $(v,w)\in\mathcal{A}$ in parallel: if the first end of the predecessor of $(v,w)$ in $\mathcal{A}$ is not $v$ then write the index of $(v,w)$ in $\mathcal{A}^{\prime}$ into $v.l$ in the private memory of (the processor corresponding to) $v$.
  3. 3.
    For each item $(v,w)\in\mathcal{A}$ in parallel: if the first end of the successor of $(v,w)$ in $\mathcal{A}$ is not $v$ then write the index of $(v,w)$ in $\mathcal{A}$ into $v.r$ in the private memory of $v$.
  4. 4.
    For each vertex $v\in V(G^{\prime})$: if $v.r-v.l\geq(\log n)^{90}$ then (approximate) compact the subarray from index $v.l$ to index $v.r$ (inclusive) in $\mathcal{A}$.
@@ 7.4.3 Computing the Edge Set in Interweave
During the execution of Connectivity$(G)$, w.p. $1-n^{-7}$, for any phase $i\in[0,10\log\log n)$ and for any vertex $v\in V(G^{\prime})$, if $\deg_{G^{\prime}}(v)\geq(\log n)^{10^{6}\cdot 1.1^{i}}$, then $v.p\notin V(G^{\prime})\backslash V(E_{\text{filter}})$ in Step 8 of Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$.
@@
W.p. $1-n^{-8}$, the edge set $E^{\prime}$ in Step 8 of Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$ in any phase $i\in[0,10\log\log n)$ of Connectivity$(G)$ satisfies $|E^{\prime}|\leq n/(\log n)^{9}$.
@@
W.p. $1-n^{-3}$, the edge set $E^{\prime}$ in Step 8 of Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$ in any phase $i\in[0,10\log\log n)$ of Connectivity$(G)$ can be computed in $O(\log b)$ time and $n/(\log n)^{7}$ work.
@@ 7.6 The Running Time and Correctness of Connectivity
During the execution of Connectivity$(G)$, w.p. $1-1/(\log n)^{2}$, for any integer $i\in[0,10\log\log n)$, phase $i$ in Step 5 takes $O(\log b)$ time unless $E_{\text{filter}}=\emptyset$, where $b=(\log n)^{100\cdot 1.1^{i}}$.
@@
During the execution of Connectivity$(G)$, w.p. $1-1/(\log n)^{7}$, for any integer $i\in[0,10\log\log n)$, if the Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$ in phase $i$ returns $\emptyset$, then the Remain$(G^{\prime},H_{1})$ in Step 3 of Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$ takes $O(\log d+\log\log n)$ time where $d$ is the maximum diameter of $G$, and Connectivity$(G)$ computes all the connected components of $G$.
@@ 8.1 Bounding the Size of Edge Set with High Probability
For all inputs $x=(x_{1},x_{2},\cdots,x_{n})\in\mathcal{X}^{n}$ where $x^{(j)}=(x_{1},\cdots,x_{j-1},x_{j+1},\cdots,x_{n})$ is the $x$ when $j^{\text{th}}$ element is dropped.
@@
Let $E^{\prime}$ be the set of edges by end of Step 8 of Interweave$(G^{\prime},H_{1},H_{2},E_{\text{filter}},i)$. Then, $|E^{\prime}|\leq 10^{9}\cdot n$ with the probability of at least $1-n^{-6}$.
@@ Appendix A Conditional Lower Bound
In this section, we show that our time bound is optimal when the input graph has component-wise spectral gap at most $1/\operatorname{poly}(\log(n))$ conditioning on the widely believed 2-Cycle Conjecture (see below). The approach is similar to that of [BDE+19], who gives an $\Omega(\log d)$ conditional lower bound, where $d\geq\operatorname{poly}(\log(n))$ is the diameter.
Conjecture A.1 (2-Cycle Conjecture [RVW18, ASW19, ASS+18, BDE+19]).
@@
For simplicity, we state the theorem in the MPC model so that our proof can be based on the MPC model. The conditional lower bound easily holds in any PRAM model with $O(m+n)$ total memory by a simple simulation [ASW19, ASS+18]: if there exists a PRAM algorithm that distinguishes between one cycle from two cycles in $o(\log n)$ time, then an MPC with $n^{1-\Omega(1)}$ space per processor can simulate this PRAM algorithm with $O(1)$ overhead, which violates the conjecture.
@@
Each round of the algorithm takes $o(\log(1/\lambda^{\prime}))$ MPC rounds w.h.p.
@@
The edge deletion at the beginning of the round and the contraction at the end of the round can be implemented in $O(1)$ MPC rounds (e.g., see [ASS+18]). It remains to show that the graph $G^{\prime}$ ($G$ after edge deletions) has spectral gap at least $\lambda^{\prime}$ w.h.p. at the beginning of each round. [^22] If graph $G^{\prime}$ has component-wise spectral gap greater than $\lambda^{\prime}$, the algorithm creates a dummy component of spectral gap $\lambda^{\prime}$ and adds it into $G^{\prime}$ such that $G^{\prime}$ has component-wise spectral gap exactly $\lambda^{\prime}$ and thus ALG is applicable on $G^{\prime}$. The dummy component has no edge with the original $G^{\prime}$ so does not influence the computation of other components. Moreover, the dummy component has at most $m$ edges (e.g. a path of fixed length) and can be created and deleted at the end of this round in $O(1)$ MPC rounds.
@@
By Lemma A.3, the total number of rounds are at most
@@ Appendix B Edge Sampling Does Not Preserve Diameter
In this section we show that edge sampling does not preserve the diameter of graphs, which partially explains why an $O(\log d)$-time, linear-work PRAM algorithm is hard to obtain.
We provide an undirected graph $G$ with a diameter of $O(\log n)$ such that if we sample every edge in $G$ with probability of $p=\frac{1}{\log n}$, the largest diameter in the connected components of $G$ would be at least $\frac{n}{\operatorname{poly}(\log n)}$ where $n$ is the number of vertices in the graph. This shows that edge sampling can dramatically increase the diameter of a graph.
@@
To show the connectivity, first consider the vertices $z_{1},\cdots,z_{(\log n)^{3}}$. We show that these vertices form a connected subgraph in $\widetilde{G}$. Notice that these vertices form a complete induced subgraph in $G$. Since we sample each edge with probability $p=1/\log n$, we can say that in $\widetilde{G}$, the subgraph induced by these vertices has the same distribution as $G((\log n)^{3},p)$. Here, $G(n,p)$ is used to denote the Erdős–Rényi graph with $n$ vertices where each edge is included in it independently with probability $p$. It is known that $G(n,p)$ is connected with probability $e^{-e^{-c}}$ if $p\geq\frac{(1+c)\ln n}{n}$ [ER+60]. Recall that in the random subgraph $G((\log n)^{3},p)$, we have $p=1/\log n$. Thus, $p\gg\frac{3\ln n\ln((\log n)^{3})}{(\log n)^{3}}$ for large enough $n$ which implies the connectivity of this random graph w.p. $1-1/n^{2}$.